OpenStudy (mini):

I need help with the MAD :) I am not good at that. 2. A wildlife refuge in South America has howler monkeys and spider monkeys. A biologist working there randomly selected 8 adults of each type of monkey, weighed them, and recorded their weights in pounds. Howler monkey: {15, 16, 17, 17, 17, 17, 18, 19} Spider monkey: {12, 13, 13, 14, 14, 14, 16, 16} (a) Calculate the mean and MAD for each type of monkey.

OpenStudy (michele_laino):

the mean values are: \[\begin{gathered} {w_H} = \frac{{15 + 16 + {\text{ }}17 + {\text{ }}17 + {\text{ }}17 + {\text{ }}17 + {\text{ }}18 + {\text{ }}19}}{8} = ...? \hfill \\ \hfill \\ {w_S} = \frac{{12 + {\text{ }}13 + {\text{ }}13 + {\text{ }}14 + {\text{ }}14 + {\text{ }}14 + {\text{ }}16 + {\text{ }}16}}{8} = ...? \hfill \\ \end{gathered} \] wherein the subscript \(H\) stands for \(howler\), and the subscript \(S\) stands for \(spider\)

OpenStudy (shy_boy):

To find the mean absolute deviation of the data, start by finding the mean of the data set. Find the sum of the data values, and divide the sum by the number of data values. Find the absolute value of the difference between each data value and the mean: |data value – mean|. Find the sum of the absolute values of the differences. Divide the sum of the absolute values of the differences by the number of data values. This is how we do it

OpenStudy (mini):

Sorry, I was writting happy birthday to someone :)

OpenStudy (mini):

Ok, well thank you guys :)

OpenStudy (michele_laino):

:)

OpenStudy (shy_boy):

yw :)